Jones covered a distance of 50 miles on his first trip. On a late trip he traveled 300 miles while going three times as fast. His new time compared with the old time was
ANSWER-
(a) three times as much
(b) twice as much
(c) the same
(d) half as much
(e) a third as much

Answer by Alan3354(69443)

Jones covered a distance of 50 miles on his first trip. On a late trip he traveled 300 miles while going three times as fast. His new time compared with the old time was
-----------
r = speed for 50 miles
------
t1 = 50/r
t2 = 300/(3r) = 100/r
==========